The rate of the elementary reaction

  • C2H2 + O2ebab9d75-6e9a-4c80-9bcd-b491b668f097.gifC2H + HO2

has been studied as a function of temperature between 300 and 2500 K. The following data were obtained for the rate constant k:

Temperature (K) k (L mol-1 s-1)

300

6.25×10-45

740

1.19×10-12

1180

1.91×10-4

1620

1.07



(a) Calculate the activation energy of this reaction.   kJ mol-1

(b) Calculate the factor A in the Arrhenius equation for the temperature dependence of the rate constant.  L mol-1 s-1
Answer #1

Givew at temperature = 740k (T) Kyuo=1.1981078 I not s-1 (ki) at Temperature = 1180 (P2) 1180 = 1.91810-4 (molt gal (ka) Tue6) [ 180 — 70) 1. 19 4 .33 х 8-39 | 1180x7чо 1.1 x 8 са чи 1. 199.3osx 3 зучение 1. 19 9. 303x8:31Ч а = 9. 303x8-319 x 118log k loga - I Ea 2.303 RI. Ay Ashenius parometter Putting the values log 1.19x1072 - log A- 311791.82 2.303x 8.3148740 loga